# Quantum theory of surface lattice resonances

### v1

@@ Abstract
The collective interactions of nanoparticles arranged in periodic structures give rise to high-$Q$ in-plane diffractive modes known as surface lattice resonances (SLRs). While these resonances and their broader implications have been extensively studied within the framework of classical electrodynamics and linear response theory, a quantum optical theory capable of describing the dynamics of these structures, especially in the presence of material nonlinearities, beyond the use of ad hoc few-mode approximations is largely missing. To this end, we consider here a lattice of metallic nanoparticles coupled to the background electromagnetic field and derive the quantum input-output relations within the electric dipole approximation. As a first application, we discuss coupling of the nanoparticle array to an external array of emitters and illustrate how the formalism can be extended to molecular optomechanics. Specifically, we show how the high $Q$-factors provided by SLRs can facilitate coupling between the nanoparticle array and collective vibrational modes. Finally, we examine a scenario in which the metallic nanoparticles themselves are replaced with saturable excitonic emitters, illustrating how the nonlinearity of the emitters can be used to switch the SLR condition between different electronic transitions. Our work provides a seamless framework to rigorously model SLR phenomena with quantum emitters without resorting to phenomenological treatments of either the electromagnetic environment or the material degrees of freedom.
@@
This paper is organized as follows. We start by deriving the optical response of an array of nanoparticles to a quantized electromagnetic field in Sec. I and illustrate how SLRs emerge due to the coupling between LSPs and diffractive modes of the lattice. This exercise provides results that are consistent with classical coupled-dipole methods, yet starting from a Hamiltonian formalism lends itself naturally for generalizations involving quantum emitters with complex internal structure. In particular, we apply the formalism towards molecular optomechanics in Sec. II and show how the high $Q$-factors exhibited by SLRs facilitate the coupling between lattice resonances and collective molecular vibrational modes. Finally, in Sec. III, we consider replacing the nanoparticles with saturable emitters such as molecules, giving rise to excitonic SLRs. We then show how the emitter nonlinearity can be harnessed to switch the SLR condition between distinct electronic states. Our Hamiltonian formalism provides a transparent pathway towards the description of complex material nonlinearities in the optical response of plasmonic arrays and is completely compatible with more sophisticated approaches to treat the plasmonic media (e.g., macroscopic QED Buhmann and Welsch (2007); Wang et al. (2019); Feist et al. (2021)).
I Model
@@
Figure 1: (a) 1D array of MNPs with lattice spacing $a$, modeled as quantum harmonic oscillators with annihilation operator $\hat{A}_{j}$, which are excited by an electric input field $\hat{\mathbf{E}}_{\mathbf{k}}^{\mathrm{in}}(\mathbf{r},t)$ incident at a wavevector $\mathbf{k}$. In-plane diffractive modes at wavevectors $\mathbf{k}_{\mathrm{RA}}$ lead to the formation of SLRs (red shaded areas indicate regions of field enhancement). (b) Extinction spectrum for an array of $M=8\cdot 10^{3}$ MNPs with a resonance wavelength of $\lambda_{0}=500\>\mathrm{nm}$, a lattice spacing of $a=550\>\mathrm{nm}$, and a single-particle radiative linewidth of $\Gamma_{0}^{\mathrm{rad}}=0.5\>\mathrm{eV}$. The dashed white lines show the dispersion of the Rayleigh anomalies (RAs) corresponding to the first diffractive order ($m=1$). The lower panel shows a cross section through the extinction profile at $|\mathbf{k}_{\parallel}|=0$ as indicated by the vertical green dashed line in the upper panel. The peak corresponds to the SLR while the dip arises from the RA. Here, the polarization of the incoming light field and the orientation of the dipoles are chosen orthogonal to the direction of the chain, i.e., $\mathbf{E}_{0}\parallel\boldsymbol{\mu}_{0}$, $\boldsymbol{\mu}_{0}\perp\mathbf{r}_{\Delta}$.
@@
Let us consider an array of $M$ MNPs located at positions $\mathbf{r}_{j}$ ($j\in\{1,\ldots,M\}$) which we model as quantum harmonic oscillators with resonance frequency $\omega_{0}$ and electric dipole operator $\hat{\boldsymbol{\mu}}_{0}^{j}=\boldsymbol{\mu}_{0}(\hat{A}_{j}^{\phantom{\dagger}}+\hat{A}_{j}^{\dagger})$. Here, the annihilation and creation operators follow the standard bosonic commutation relation $[\hat{A}_{j}^{\phantom{\dagger}},\hat{A}_{j^{\prime}}^{\dagger}]=\delta_{j,j^{\prime}}$. We consider a regular lattice of particles spaced between each other by a distance $a$, as illustrated in Fig. 1(a). While the theoretical treatment employed in the following holds for both 1D and 2D lattices, for the sake of simplicity we will focus on the 1D case. The lattice is assumed to be quasi-infinite which allows us to impose periodic boundary conditions. The array of nanoparticles is then described by the Hamiltonian
^$\hat{\mathcal{H}}_{\text{MNP}}=\hbar\sum_{j=1}^{M}\omega_{0}^{\phantom{\dagger}}\hat{A}_{j}^{\dagger}\hat{A}_{j}^{\phantom{\dagger}}.$ (1)
@@
^$\hat{\mathcal{H}}_{\text{vac}}=\hbar\sum_{\mathbf{k},\lambda}\omega_{\mathbf{k}}^{\phantom{\dagger}}\hat{a}_{\mathbf{k},\lambda}^{\dagger}\hat{a}_{\mathbf{k},\lambda}^{\phantom{\dagger}},$ (2)
[…]
^$\hat{\mathcal{H}}_{\text{int}}=\hbar\sum_{j=1}^{M}\sum_{\mathbf{k},\lambda}\left(g_{\mathbf{k},\lambda}^{\phantom{\dagger}}\hat{A}_{j}^{\dagger}\hat{a}_{\mathbf{k},\lambda}^{\phantom{\dagger}}\mathrm{e}^{\mathrm{i}\mathbf{k}\cdot\mathbf{r}_{j}}+\mathrm{H.c.}\right),$ (3)
@@
Our objective is now to find a description for the response of the MNP array to a plane-wave input field $\hat{\mathbf{E}}_{\mathbf{k}}^{\mathrm{in}}(\mathbf{r},t)=-\mathrm{i}\mathbf{E}_{0}(\mathrm{e}^{-\mathrm{i}\mathbf{k}\cdot\mathbf{r}}\hat{a}_{\mathbf{k},\lambda}^{\mathrm{in}}(t)-\mathrm{H.c.})$, with polarization vector $\mathbf{E}_{0}$. This can be achieved by deriving the equations of motion for the dipole operators as well as for the the electric field amplitudes in the Heisenberg picture, followed by an elimination of the electromagnetic degrees of freedom, to obtain a reduced description for the matter part only [see Supplemental Material (SM) for detailed steps of the derivation]. This procedure is analogous to the derivation of generalized Langevin equations Zwanzig (2001) and closely resembles the treatment of collective emission of atomic dipoles in quantum optics Dicke (1954); Lehmberg (1970a, b); Gardiner and Zoller (2000). Crucially, however, it does not invoke the Markovian assumption typically applied to such systems, thereby incorporating all retardation effects which are essential for obtaining the narrow lineshape associated with SLRs. Given the harmonic nature of ^$\hat{\mathcal{H}}_{\text{MNP}}$, ^$\hat{\mathcal{H}}_{\text{vac}}$, and ^$\hat{\mathcal{H}}_{\text{int}}$, we afford an analytical expression for the dipole amplitudes in Fourier domain
@@
Figure 2: Intensity profile $|\mathbf{E}(\mathbf{r})|^{2}$ radiated by a 1D chain of MNPs (indicated as silver dots) with a resonance wavelength of $\lambda_{0}=500\>\mathrm{nm}$ and a radiative linewidth of $\Gamma_{0}^{\mathrm{rad}}=0.5\>\mathrm{eV}$, for a lattice spacing of (a) $a=300\>\mathrm{nm}$, and (b) $a=550\>\mathrm{nm}$. The field distribution is plotted in the $xz$-plane and we have assumed normally-incident illumination of the array ($\theta_{\mathrm{inc}}=0$) with a polarization vector along the $y$ direction, i.e., $\mathbf{E}_{0}\parallel\boldsymbol{\mu}_{0}\parallel\mathbf{e}_{y}$. In (a), the electric field is evaluated at the frequency corresponding to the original LSP resonance ($\omega_{0}=2\pi c/\lambda_{0}$), while in (b) it is evaluated at the red-shifted SLR frequency.
This is in full agreement with the result obtained from the classical electromagnetical description (see, e.g., Eq. (10) of Ref. Cherqui et al. (2019)). More generally, the effective polarizability tensor of the MNP lattice can always be related to the single-particle polarizability $\boldsymbol{\alpha}_{0}(\omega)$ via $[\boldsymbol{\alpha}_{\mathbf{q}}^{\mathrm{eff}}]^{-1}(\omega)=\boldsymbol{\alpha}_{0}^{-1}(\omega)-\mathcal{S}_{\mathbf{q}}(\omega)/(\boldsymbol{\mu}_{0}\cdot\boldsymbol{\mu}_{0}^{\top})$. The extinction spectrum (describing the combined effects of absorption and scattering) of the nanoparticle array is defined as Cherqui et al. (2019)
@@
II Application 1: Molecular optomechanics with SLRs
@@
The field of optomechanics is concerned with the interaction between light and mechanical motion. In a prototypical optomechanical setup, circulating photons exert a force on the mirrors due to radiation pressure Aspelmeyer et al. (2014). The mutual interplay between light and motion gives rise to a variety of applications of optomechanical devices, e.g., as sensors Krause et al. (2012); Fogliano et al. (2021) or frequency-converters Hill et al. (2012). Interestingly, an analogy can be found between the standard optomechanical setup and surface-enhanced Raman scattering (SERS) where the plasmonic field takes the role of the cavity mode and the molecular vibration takes that of the mechanical oscillator Roelli et al. (2016). While SERS is typically understood as a classical process, parametric amplification phenomena experimentally observed in these systems sparked interest in recent years to translate concepts from quantum cavity optomechanics Chen et al. (2021), leading to a quantum description of the interaction between molecular vibrations and plasmonic fields Neuman et al. (2019); Patra et al. (2023). Given that molecular optomechanics often suffers from lossy plasmonic modes, it is worthwhile exploring whether the high $Q$-factors offered by SLRs can provide an advantage. In particular, this may help achieve the desired resolved sideband regime in optomechanics, which is characterized by vibrational frequencies exceeding the optical mode linewidth Zou and Schatz (2005), a crucial requirement for selectively addressing individual Stokes/anti-Stokes sidebands.
To this end, we now consider an additional array of $M$ (identical) molecules at positions $\mathbf{r}_{j}^{\mathrm{m}}$ which are displaced by some fixed vector $\mathbf{r}_{\mathrm{m}}$ from the MNP lattice, i.e., $\mathbf{r}_{j}^{\mathrm{m}}=\mathbf{r}_{j}+\mathbf{r}_{\mathrm{m}}$. In the following we will aim at positioning the molecules into the SLR ‘hotspots’, as illustrated in Fig. 3(a). The general optomechanical Hamiltonian expresses as
^$\hat{\mathcal{H}}_{\text{OM}}=-\sum_{j}\hat{\mathbf{p}}_{j}^{R}\cdot\hat{\mathbf{E}}(\mathbf{r}_{j}^{\mathrm{m}}),$ (8)
@@
with the induced Raman dipole of the molecule $\hat{\mathbf{p}}_{j}^{R}=\alpha_{\mathrm{m}}(Q_{j})\hat{\mathbf{E}}(\mathbf{r}_{j}^{\mathrm{m}})$, where we consider the (scalar) molecular polarizability $\alpha_{\mathrm{m}}(\hat{Q}_{j})$ to depend on a single quantized, harmonic nuclear coordinate $\hat{Q}_{j}=Q_{\mathrm{zpm}}(\hat{b}_{j}^{\dagger}+\hat{b}_{j})$. Here, the zero-point motion is given by $Q_{\mathrm{zpm}}=\sqrt{\hbar/(2m_{\mathrm{vib}}\omega_{\mathrm{vib}})}$ with $\omega_{\mathrm{vib}}$ the vibrational frequency and $m_{\mathrm{vib}}$ the reduced mass of the vibrational mode. The free Hamiltonian of the nuclear degrees of freedom is then simply given by ^$\hat{\mathcal{H}}_{\mathrm{vib}}=\sum_{j}\omega_{\mathrm{vib}}\hat{b}_{j}^{\dagger}\hat{b}_{j}^{\phantom{\dagger}}$. Taylor expansion of the polarizability with respect to the nuclear coordinates around the equilibrium configuration $Q_{0}$ then gives rise to a coupling between the electromagnetic field and the nuclear motion
@@
^$\hat{\mathcal{H}}_{\mathrm{OM}}=\hbar\sum_{j=1}^{M}\sum_{\mathbf{k},\lambda}\left[g_{\mathbf{k},\lambda}^{\mathrm{OM}}\hat{a}_{\mathbf{k},\lambda}\mathrm{e}^{\mathrm{i}\mathbf{k}\cdot\mathbf{r}_{j}^{\mathrm{m}}}\mathrm{e}^{\mathrm{i}\omega_{\ell}t}+\mathrm{H.c.}\right](\hat{b}_{j}^{\dagger}+\hat{b}_{j}^{\phantom{\dagger}}),$ (11)
@@
$\mathcal{S}_{\mathbf{q}}^{\mathrm{OM}}(\omega)\!=\frac{\mu_{0}p\omega^{2}}{c^{2}\hbar\epsilon_{0}}\sum_{\Lambda}^{M-1}\mathrm{e}^{-\mathrm{i}\mathbf{q}\cdot(\mathbf{r}_{\Lambda}+\mathbf{r}_{\mathrm{m}})}\boldsymbol{\varepsilon}_{\boldsymbol{\mu}}\!\cdot\!\mathbf{G}(\mathbf{r}_{\Lambda}\!+\mathbf{r}_{\mathrm{m}},\omega)\!\cdot\!\boldsymbol{\varepsilon}_{\mathbf{p}}.$ (14)
@@
It might be possible, however, that under specific design conditions—for example, if a suitable SLR is engineered such that its broad background resides at lower energies than $\omega_{0}$—the blue-detuned scenario becomes more favorable. Verifying such a configuration requires going beyond the RWA and keeping both rotating and counter-rotating terms in the optomechanical Hamiltonian (see SM).
@@
The resulting extinction spectra of the nanoparticle array are plotted in Figs. 3(c) and (d) in the red- and blue-detuned regimes. Here, we placed the molecules in the regions of field enhancement between the MNPs, i.e., at $\mathbf{r}_{\mathrm{m}}=0.5a\mathbf{e}_{x}$ and assumed dipole orientations $\mathbf{p}\!\parallel\!\boldsymbol{\mu}_{0}\perp\!\mathbf{r}_{\Delta}$. The red and blue sideband regime have distinct optical signatures due to the different underlying exchange processes. In the anti-Stokes regime, where the laser is detuned below the SLR resonance (red-detuned) by the molecular vibrational frequency, the spectrum exhibits a dip in extinction corresponding to optomechanically-induced transparency (OMIT), or in the strong coupling regime, a normal mode splitting due to coherent energy exchanges between the collective lattice resonance and the vibrational mode Aspelmeyer et al. (2014). In contrast, the blue sideband regime, with the pump laser detuned above the SLR resonance, shows a broadened peak in extinction and does not exhibit strong coupling.

### v2

@@ Abstract
The collective interactions of nanoparticles arranged in periodic structures give rise to high-$Q$ in-plane diffractive modes known as surface lattice resonances (SLRs). While these resonances and their broader implications have been extensively studied within the framework of classical electrodynamics and linear response theory, a quantum optical theory capable of describing the dynamics of these structures, especially in the presence of material nonlinearities, beyond the use of ad hoc few-mode approximations is largely missing. To this end, we consider here a lattice of metallic nanoparticles coupled to the background electromagnetic field and derive the quantum input-output relations within the electric dipole approximation. As a first application, we discuss coupling of the nanoparticle array to an external array of emitters and illustrate how the formalism can be extended to molecular optomechanics. Specifically, we show how the high $Q$-factors provided by SLRs can facilitate coupling between the nanoparticle array and collective vibrational modes. Finally, we examine a scenario in which the metallic nanoparticles themselves are replaced with saturable excitonic emitters, illustrating how the nonlinearity of the emitters can be used to switch the SLR condition between different electronic transitions. We further show, using a perturbative approach that explicitly accounts for population dynamics, how the resulting switching dynamics can be observed and characterized in pump–probe experiments. Our work provides a seamless framework to rigorously model SLR phenomena with quantum emitters without resorting to phenomenological treatments of either the electromagnetic environment or the material degrees of freedom.
I Motivation
@@
This paper is organized as follows. We start by deriving the optical response of an array of nanoparticles to a quantized electromagnetic field and illustrate how SLRs emerge due to the coupling between LSPs and diffractive modes of the lattice. This exercise provides results that are consistent with classical coupled-dipole methods, yet starting from a Hamiltonian formalism lends itself naturally for generalizations involving quantum emitters with complex internal structure. In particular, we apply the formalism towards molecular optomechanics and show how the high $Q$-factors exhibited by SLRs facilitate the coupling between lattice resonances and collective molecular vibrational modes. Finally, we consider replacing the nanoparticles with saturable emitters such as molecules, giving rise to excitonic SLRs. We then show how the emitter nonlinearity can be harnessed to switch the SLR condition between distinct electronic states under the assumption of static populations. We then extend the treatment to include dynamical population effects in a perturbative fashion, demonstrating that the resulting switching dynamics can be observed and characterized in pump–probe experiments. Our Hamiltonian formalism provides a transparent pathway towards the description of complex material nonlinearities in the optical response of plasmonic arrays and is completely compatible with more sophisticated approaches to treat the plasmonic media (e.g., macroscopic QED Buhmann and Welsch (2007); Wang et al. (2019); Feist et al. (2021)).
II Model
@@
Figure 1: (a) 1D array of MNPs with lattice spacing $a$, modeled as quantum harmonic oscillators with annihilation operator $\hat{A}_{j}$, which are excited by an electric input field $\hat{\mathbf{E}}_{\mathbf{k}}^{\mathrm{in}}(\mathbf{r},t)$ incident at a wavevector $\mathbf{k}$. In-plane diffractive modes at wavevectors $\mathbf{k}_{\mathrm{RA}}$ lead to the formation of SLRs (red shaded areas indicate regions of field enhancement). (b) Extinction spectrum for an array of $M=8\cdot 10^{3}$ MNPs with a resonance wavelength of $\lambda_{0}=500\>\mathrm{nm}$, a lattice spacing of $a=550\>\mathrm{nm}$, and a single-particle radiative linewidth of $\Gamma_{0}^{\mathrm{rad}}=0.5\>\mathrm{eV}$. The dashed white lines show the dispersion of the Rayleigh anomalies (RAs) corresponding to the first diffractive order ($m=1$). The lower panel shows a cross section through the extinction profile at $|\mathbf{k}_{\parallel}|=0$ as indicated by the vertical green dashed line in the upper panel. The peak corresponds to the SLR while the dip arises from the RA. Here, the polarization of the incoming light field and the orientation of the dipoles are chosen orthogonal to the direction of the chain, i.e., $\mathbf{E}_{0}\parallel\boldsymbol{\mu}_{0}$, $\boldsymbol{\mu}_{0}\perp\mathbf{r}_{\Lambda}$.
@@
Let us consider an array of $M$ MNPs located at positions $\mathbf{r}_{j}$ ($j\in\{1,\ldots,M\}$) which we model as quantum harmonic oscillators with resonance frequency $\omega_{0}$ and electric dipole operator $\hat{\boldsymbol{\mu}}_{0}^{j}=\boldsymbol{\mu}_{0}(\hat{A}_{j}^{\phantom{\dagger}}+\hat{A}_{j}^{\dagger})$. Here, the annihilation and creation operators follow the standard bosonic commutation relation $\hat{A}_{j}^{\phantom{\dagger}},\hat{A}_{j^{\prime}}^{\dagger}]=\delta_{j,j^{\prime}}$. We consider a regular lattice of particles spaced between each other by a distance $a$, as illustrated in Fig. 1(a). While the theoretical treatment employed in the following holds for both 1D and 2D lattices, for the sake of simplicity we will focus on the 1D case. We note, however, that the physics of surface lattice resonances can exhibit substantial distinctions between 1D and 2D geometries due to differences in dipole coupling structure, mode multiplicity and convergence, as discussed for instance in Refs. Zou and Schatz (2004); Cherqui et al. (2019). The lattice is assumed to be quasi-infinite which allows us to impose periodic boundary conditions. The array of nanoparticles is then described by the Hamiltonian
$\hat{\mathcal{H}}_{\text{MNP}}=\hbar\sum_{j=1}^{M}\omega_{0}^{\phantom{\dagger}}\hat{A}_{j}^{\dagger}\hat{A}_{j}^{\phantom{\dagger}}.$ (1)
@@
$\hat{\mathcal{H}}_{\text{vac}}=\hbar\sum_{\mathbf{k},\lambda}\omega_{\mathbf{k}}^{\phantom{\dagger}}\hat{a}_{\mathbf{k},\lambda}^{\dagger}\hat{a}_{\mathbf{k},\lambda}^{\phantom{\dagger}},$ (2)
where $\hat{a}_{\mathbf{k},\lambda}^{\phantom{\dagger}}$ ($\hat{a}_{\mathbf{k},\lambda}^{\dagger}$) is the photon annihilation (creation) operator for a given electromagnetic field mode, $\omega_{\mathbf{k}}=c|\mathbf{k}|$ is the vacuum photonic dispersion, and the sum goes over all the wavevectors $\mathbf{k}$ which can be decomposed into in-plane and out-of-plane components $\mathbf{k}=(\mathbf{k}_{\parallel},k_{\perp})^{\top}$, as well as over the two orthogonal polarizations associated with each $\mathbf{k}$ mode which are denoted by the index $\lambda$. The interaction between the MNPs and the bath of electromagnetic modes is described within the rotating wave approximation by
$\hat{\mathcal{H}}_{\text{int}}=\hbar\sum_{j=1}^{M}\sum_{\mathbf{k},\lambda}\left(g_{\mathbf{k},\lambda}^{\phantom{\dagger}}\hat{A}_{j}^{\dagger}\hat{a}_{\mathbf{k},\lambda}^{\phantom{\dagger}}\mathrm{e}^{\mathrm{i}\mathbf{k}\cdot\mathbf{r}_{j}}+\mathrm{H.c.}\right),$ (3)
@@
Our objective is now to find a description for the response of the MNP array to a plane-wave input field $\hat{\mathbf{E}}_{\mathbf{k}}^{\mathrm{in}}(\mathbf{r},t)=-\mathrm{i}\mathbf{E}_{0}(\mathrm{e}^{-\mathrm{i}\mathbf{k}\cdot\mathbf{r}}\hat{a}_{\mathbf{k},\lambda}^{\mathrm{in}}(t)-\mathrm{H.c.})$, with polarization vector $\mathbf{E}_{0}$. This can be achieved by deriving the equations of motion for the dipole operators as well as for the the electric field amplitudes in the Heisenberg picture, followed by an elimination of the electromagnetic degrees of freedom, to obtain a reduced description for the matter part only [see Supplemental Material (SM) for detailed steps of the derivation]. This procedure is analogous to the derivation of generalized Langevin equations Zwanzig (2001) and closely resembles the treatment of collective emission of atomic dipoles in quantum optics Dicke (1954); Lehmberg (1970a, b); Gardiner and Zoller (2000). Crucially, however, it does not invoke the Markovian assumption typically applied to such systems, thereby incorporating all retardation effects which are essential for obtaining the narrow lineshape associated with SLRs. Given the harmonic nature of $\hat{\mathcal{H}}_{\text{MNP}}$, $\hat{\mathcal{H}}_{\text{vac}}$, and $\hat{\mathcal{H}}_{\text{int}}$, we afford an analytical expression for the dipole amplitudes in Fourier domain
@@
Figure 2: Intensity profile $|\mathbf{E}(\mathbf{r})|^{2}$ radiated by a 1D chain of MNPs (indicated as silver dots) with a resonance wavelength of $\lambda_{0}=500\>\mathrm{nm}$ and a radiative linewidth of $\Gamma_{0}^{\mathrm{rad}}=0.5\>\mathrm{eV}$, for a lattice spacing of (a) $a=300\>\mathrm{nm}$, and (b) $a=550\>\mathrm{nm}$. The field distribution is plotted in the xzxz-plane and we have assumed normally-incident illumination of the array ($\theta_{\mathrm{inc}}=0$) with a polarization vector along the $y$ direction, i.e., $\mathbf{E}_{0}\parallel\boldsymbol{\mu}_{0}\parallel\mathbf{e}_{y}$. In (a), the electric field is evaluated at the frequency corresponding to the original LSP resonance ($\omega_{0}=2\pi c/\lambda_{0}$), while in (b) it is evaluated at the red-shifted SLR frequency.
This is in full agreement with the result obtained from the classical electromagnetical description (see, e.g., Eq. (10) of Ref. Cherqui et al. (2019)). More generally, the effective polarizability tensor of the MNP lattice can always be related to the single-particle polarizability $\boldsymbol{\alpha}_{0}(\omega)$ via $\boldsymbol{\alpha}_{\mathbf{q}}^{\mathrm{eff}}]^{-1}(\omega)=\boldsymbol{\alpha}_{0}^{-1}(\omega)-\mathcal{S}_{\mathbf{q}}(\omega)/(\boldsymbol{\mu}_{0}\cdot\boldsymbol{\mu}_{0}^{\top})$. The extinction spectrum (describing the combined effects of absorption and scattering) of the nanoparticle array is defined as Cherqui et al. (2019)
@@
III Application 1: Molecular optomechanics with SLRs
@@
The field of optomechanics is concerned with the interaction between light and mechanical motion. In a prototypical optomechanical setup, circulating photons exert a force on the mirrors due to radiation pressure Aspelmeyer et al. (2014). The mutual interplay between light and motion gives rise to a variety of applications of optomechanical devices, e.g., as sensors Krause et al. (2012); Fogliano et al. (2021) or frequency-converters Hill et al. (2012). Interestingly, an analogy can be found between the standard optomechanical setup and surface-enhanced Raman scattering (SERS) where the plasmonic field takes the role of the cavity mode and the molecular vibration takes that of the mechanical oscillator Roelli et al. (2016). While SERS is typically understood as a classical process, parametric amplification phenomena experimentally observed in these systems sparked interest in recent years to translate concepts from quantum cavity optomechanics Chen et al. (2021), leading to a quantum description of the interaction between molecular vibrations and plasmonic fields Neuman et al. (2019); Patra et al. (2023). Given that molecular optomechanics often suffers from lossy plasmonic modes, it is worthwhile exploring whether the high $Q$-factors offered by SLRs can provide an advantage. In particular, this may help achieve the desired resolved sideband regime in optomechanics, which is characterized by vibrational frequencies exceeding the optical mode linewidth Zou and Schatz (2005), a crucial requirement for selectively addressing individual Stokes/anti-Stokes sidebands. SLRs have e.g. been experimentally studied for SERS signal enhancement in Refs. Alam et al. (2020); Khinevich et al. ; Andrzejewska et al. (2025) . Although here, we focus on molecular vibrations, it is also noteworthy that optomechanical oscillations can occur directly within the nanoparticles themselves which has been observed via time-resolved spectroscopy and does not require the coupling to external emitters Hartland (2011); Juodėnas et al. (2020).
To this end, we now consider an additional array of $M$ (identical) molecules at positions $\mathbf{r}_{j}^{\mathrm{m}}$ which are displaced by some fixed vector $\mathbf{r}_{\mathrm{m}}$ from the MNP lattice, i.e., $\mathbf{r}_{j}^{\mathrm{m}}=\mathbf{r}_{j}+\mathbf{r}_{\mathrm{m}}$. While such a perfectly ordered molecular array may be challenging to realize experimentally, we adopt this configuration here because it preserves translational symmetry and enables momentum conservation, thereby greatly simplifying the theoretical analysis. Further, we focus on SLR modes with wave vectors much larger than the inverse intermolecular spacing, so that we expect disorder-induced variations to effectively average out over these longer wavelengths. However, recent advances in nanoprinting techniques have demonstrated the ability to position organic molecules with high precision Hail et al. (2019), offering promising routes toward realizing such ordered hybrid architectures. In the following we will aim at positioning the molecules into the SLR ‘hotspots’, as illustrated in Fig. 3(a). The general optomechanical Hamiltonian expresses as
$\hat{\mathcal{H}}_{\text{OM}}=-\sum_{j}\hat{\mathbf{p}}_{j}^{R}\cdot\hat{\mathbf{E}}(\mathbf{r}_{j}^{\mathrm{m}}),$ (8)
@@
with the induced Raman dipole of the molecule $\hat{\mathbf{p}}_{j}^{R}=\alpha_{\mathrm{m}}(Q_{j})\hat{\mathbf{E}}(\mathbf{r}_{j}^{\mathrm{m}})$, where we consider the (scalar) molecular polarizability $\alpha_{\mathrm{m}}(\hat{Q}_{j})$ to depend on a single quantized, harmonic nuclear coordinate $\hat{Q}_{j}=Q_{\mathrm{zpm}}(\hat{b}_{j}^{\dagger}+\hat{b}_{j})$. Here, the zero-point motion is given by $Q_{\mathrm{zpm}}=\sqrt{\hbar/(2m_{\mathrm{vib}}\omega_{\mathrm{vib}})}$ with $\omega_{\mathrm{vib}}$ the vibrational frequency and $m_{\mathrm{vib}}$ the reduced mass of the vibrational mode. The free Hamiltonian of the nuclear degrees of freedom is then simply given by $\hat{\mathcal{H}}_{\mathrm{vib}}=\sum_{j}\omega_{\mathrm{vib}}\hat{b}_{j}^{\dagger}\hat{b}_{j}^{\phantom{\dagger}}$. Taylor expansion of the polarizability with respect to the nuclear coordinates around the equilibrium configuration $Q_{0}$ then gives rise to a coupling between the electromagnetic field and the nuclear motion
@@
$\hat{\mathcal{H}}_{\mathrm{OM}}=\hbar\sum_{j=1}^{M}\sum_{\mathbf{k},\lambda}\left[g_{\mathbf{k},\lambda}^{\mathrm{OM}}\hat{a}_{\mathbf{k},\lambda}\mathrm{e}^{\mathrm{i}\mathbf{k}\cdot\mathbf{r}_{j}^{\mathrm{m}}}\mathrm{e}^{\mathrm{i}\omega_{\ell}t}+\mathrm{H.c.}\right](\hat{b}_{j}^{\dagger}+\hat{b}_{j}^{\phantom{\dagger}}),$ (11)
@@
$\mathcal{S}_{\mathbf{q}}^{\mathrm{OM}}(\omega)\!=\frac{\mu_{0}p\omega^{2}}{c^{2}\hbar\epsilon_{0}}\sum_{\Lambda}\mathrm{e}^{-\mathrm{i}\mathbf{q}\cdot(\mathbf{r}_{\Lambda}+\mathbf{r}_{\mathrm{m}})}\boldsymbol{\varepsilon}_{\boldsymbol{\mu}}\!\cdot\!\mathbf{G}(\mathbf{r}_{\Lambda}\!+\mathbf{r}_{\mathrm{m}},\omega)\!\cdot\!\boldsymbol{\varepsilon}_{\mathbf{p}}.$ (14)
@@
It might be possible, however, that under specific design conditions—for example, if a suitable SLR is engineered such that its broad background resides at lower energies than $\omega_{0}$ —the blue-detuned scenario becomes more favorable. For related photonic lattices, this can, e.g., be achieved by tuning the fill factor, which modifies the band structure and enables control over the relative position of leaky and bound-state modes Lee and Magnusson (2019).
@@
The resulting extinction spectra of the nanoparticle array are plotted in Figs. 3(c) and (d) in the red- and blue-detuned regimes. Here, we placed the molecules in the regions of field enhancement between the MNPs, i.e., at $\mathbf{r}_{\mathrm{m}}=0.5a\mathbf{e}_{x}$ and assumed dipole orientations $\mathbf{p}\!\parallel\!\boldsymbol{\mu}_{0}\perp\!\mathbf{r}_{\Lambda}$. The red and blue sideband regime have distinct optical signatures due to the different underlying exchange processes. In the anti-Stokes regime, where the laser is detuned below the SLR resonance (red-detuned) by the molecular vibrational frequency, the spectrum exhibits a dip in extinction corresponding to optomechanically-induced transparency (OMIT), or in the strong coupling regime, a normal mode splitting due to coherent energy exchanges between the collective lattice resonance and the vibrational mode Aspelmeyer et al. (2014). In contrast, the blue sideband regime, with the pump laser detuned above the SLR resonance, shows a broadened peak in extinction and does not exhibit strong coupling.